# From Friendship Networks to Classroom Dynamics: Leveraging Neural Networks, Instrumental Variable and Genetic Algorithms for Optimal Educational Outcomes

### v4

@@ Appendix B Friendship questionnaire
Appendix C Examples
Example C.1.
A typical mainstream problem that fits current literature and our research context would be: if a new student joins a classroom, given the current network structure in the classroom, all classmates’ characteristics, and the new student’s characteristics, who is most likely to befriend the new student (i.e. predicting future link formation). This is a factual prediction problem as we can observe the friendship decisions made by the new student later on. Our work solves a different problem. In our work, we are interested in the following problem: given the current network structures of some 8th-grade classrooms and all students’ 6th-grade characteristics, how will the network structures change if we assign students differently two years ago at the start of these students’ 7th-grade. This is a counterfactual prediction problem. We will never observe the counterfactual scenario that we are trying to predict.
In the former question, one may rely more on graph-based measures such as node degree to make a prediction; in the latter, we use content-based measures exclusively to predict friendship. See Daud et al. (2020) for details on graph-based and content-based measures. Since the latter type of questions including our research question can only use content-based measures, there is arguably a greater need for learning about the underlying friendship preference function than the former type of questions.
More importantly, the former question is a factual prediction problem, whereas the latter is a counterfactual prediction problem. While researchers can observe how a new student forms friendships in a classroom, they never observe the same group of students being assigned to different classrooms twice and observe two sets of networks independently. Therefore, many of the techniques that we use for prediction (e.g. any method that uses existing network structure such as using GNN to aggregate neighbor information) cannot be applied or have to be applied differently to the latter question.
With example C.1, we argue that a micro-founded approach supported by economic utility theory is particularly suitable for our research question.
Example C.2.
@@ Appendix C Examples
Example C.3.
We compute the outer product of $\delta$ and $\sigma$ from example C.2 to compute the linearized propensity score $\Upsilon_{ij}$.
@@
Example C.4.
Students may choose to make friends with those of similar IQ as themselves. Higher IQ also has a positive impact on students’ 8th-grade cognitive ability test scores. We do not observe IQ and hence it is part of the error term. Then, if we run regression based on Eq(3), we may overestimate peer effect $\beta$ since we inevitably attribute some of the impact of IQ to peer effect.
Example C.5.
We continue with example C.4 where IQ is an endogenous variable.
Exclusion constraint requires the instrument (average 6th-grade class rank for all classmates) to be uncorrelated with the unobserved confounder IQ. This requirement is trivially satisfied since all classroom assignments are random conditional on school choice which is a covariate included in Eq(3).
@@
Example C.6.
A principal needs to divide 100 students into two classrooms, each consisting of 50 students. There are $\frac{100!}{50!}$ ways to assign students to classrooms.
Appendix D Explanation for the loss function
D.1 Fitted value’s MSE
@@ D.1.3 Decomposition of MSE
D.1.4 Closed-form expressions for ${\rm Bias}^{2}$ and ${\rm Var}$
@@ D.2 Downweighing variance
D.3 Using fitted $\Omega$ to select $(\mu,\kappa,\lambda)$
@@
For example, as mentioned in Appendix D.2, $(\mu,\kappa,\lambda)=(1,0,0)$ is unreasonable as the $\Omega$ matrix generated by this combination of tuning parameters violates the dense nature of classroom network.
@@
We pick $(\mu,\kappa,\lambda)=(0.2,0.3,0.3)$. The choice of hyperparameters might seem ad hoc; however, one can formalize measures of gender homophily, node centrality, and clustering to optimize these hyperparameter choices more systematically. Such a tuning method is time-efficient because we only need to fit the model once for a particular choice of $(\mu,\kappa,\lambda)$ in comparison to $k$-fold cross-validation which requires $k$ instances of the fitting.
Appendix E Linear-in-means model
@@
Linear-in-means model implicitly assumes that all students exert the same amount of peer influence on all their classmates as evidenced by the $W_{cs}$ matrix. The structural equation that contains the parameter of interest $\beta$ (i.e. Equation (3)) replaces $W_{cs}$ with $\Omega_{cs}$. The microfoundation for specification of Equation (3) is the following.
E.1 Microfoundation of Eq(3)
@@
As illustrated by example C.4, ${\rm Cov}(\Omega_{cs},\epsilon_{cs})\neq 0$. In economics, this problem is called endogeneity. We address endogeneity by constructing an instrumental variable which happens to be $W_{cs}z_{cs}$.
E.2 Comparison between linear-in-means and friendship-weighted specifications
The two $\beta$s in Eq(2) and Eq(3) have different interpretations. For linear-in-means model, $\beta$ in Eq(2) is the marginal effect of middle school classmates’ average 6th-grade class quantile on the outcome; for friendship-weighted specification, $\beta$ in Eq(3) is the marginal effect of middle school friends’ weighted average 6th-grade class quantile on the outcome. The former has very little policy implication on the average peer effect since the linear-in-means model essentially restricts how much total peer effect each student can exert on other classmates. Our friendship-weighted specification relaxes this restriction: a popular student is allowed to exert more influence on all his classmates than another less popular student. Students may also have different levels of influence on different subgroups, for example, in our empirical results, boys influence boys more and girls influence girls more. This allows us to devise average-maximizing class assignments.
Appendix F Genetic algorithm used to optimize class assignment policy
@@
     end if
end procedure
Algorithm 3 swap function
@@
end procedure
Appendix G Friendship formation pattern
We show that the friendship formation patterns that we describe in our main text are the same throughout different classroom networks with Figure LABEL:heat_maps_FIM_69_and_138. First, the gender homophily effect is the dominant predictor for friendship formation. Boys almost exclusively make friends with boys and girls make friends with girls. Second, the centrality pattern is clear in the heat maps. Some individual students are extremely influential to other students. Third, students have varying degrees of influence on other students in the same classroom. Even within the same gender, one student is considered by other students as best friends with different probabilities.
@@ Appendix H Class assignment results
We show results for multiple schools. From all figures, GA always severely penalizes a few students in pursuit of the highest average peer effect. In contrast, AFGA generates much more fair class assignment policies. In some cases, readers might be able to visually discern that the average peer effect of AFGA is higher than raw assignment as AFGA tends to have lighter color plots.

### v5

@@ Appendix A Friendship questionnaire
Appendix B Examples
Example B.1.
@@ Appendix B Examples
Example B.2.
We compute the outer product of $\delta$ and $\sigma$ from example B.1 to compute the linearized propensity score $\Upsilon_{ij}$.
@@
Example B.3.
Students may choose to make friends with those of similar IQ as themselves. Higher IQ also has a positive impact on students’ 8th-grade cognitive ability test scores. We do not observe IQ and hence it is part of the error term. Then, if we run regression based on (3.3), we may overestimate peer effect $\beta$ since we inevitably attribute some of the impact of IQ to peer effect.
Example B.4.
We continue with example B.3 where IQ is an endogenous variable.
Exclusion constraint requires the instrument (average 6th-grade class rank for all classmates) to be uncorrelated with the unobserved confounder IQ. This requirement is trivially satisfied since all classroom assignments are random conditional on school choice which is a covariate included in (3.3).
@@
Appendix C Explanation for the loss function
C.1 Fitted value’s MSE
@@ C.1.3 Decomposition of MSE
C.1.4 Closed-form expressions for ${\rm Bias}^{2}$ and ${\rm Var}$
@@ C.2 Downweighing variance
C.3 Using fitted $\Omega$ to select $(\mu,\kappa,\lambda)$
@@
For example, as mentioned in Appendix C.2, $(\mu,\kappa,\lambda)=(1,0,0)$ is unreasonable as the $\Omega$ matrix generated by this combination of tuning parameters violates the dense nature of classroom network.
@@
We pick $(\mu,\kappa,\lambda)=(0.2,3,3)$. The choice of hyperparameters might seem ad hoc; however, one can formalize measures of gender homophily, node centrality, and clustering to optimize these hyperparameter choices more systematically. Such a tuning method is time-efficient because we only need to fit the model once for a particular choice of $(\mu,\kappa,\lambda)$ in comparison to $k$-fold cross-validation which requires $k$ instances of the fitting.
Appendix D Linear-in-means model
@@
Linear-in-means model implicitly assumes that all students exert the same amount of peer influence on all their classmates as evidenced by the $W_{cs}$ matrix. The structural equation that contains the parameter of interest $\beta$ (i.e. uation (3.3)) replaces $W_{cs}$ with $\Omega_{cs}$. The microfoundation for specification of uation (3.3) is the following.
D.1 Microfoundation of (3.3)
@@
As illustrated by example B.3, ${\rm Cov}(\Omega_{cs},\epsilon_{cs})\neq 0$. In economics, this problem is called endogeneity. We address endogeneity by constructing an instrumental variable which happens to be $W_{cs}z_{cs}$.
D.2 Comparison between linear-in-means and friendship-weighted specifications
The two $\beta$s in (3.2) and (3.3) have different interpretations. For linear-in-means model, $\beta$ in (3.2) is the marginal effect of middle school classmates’ average 6th-grade class quantile on the outcome; for friendship-weighted specification, $\beta$ in (3.3) is the marginal effect of middle school friends’ weighted average 6th-grade class quantile on the outcome. The former has very little policy implication on the average peer effect since the linear-in-means model essentially restricts how much total peer effect each student can exert on other classmates. Our friendship-weighted specification relaxes this restriction: a popular student is allowed to exert more influence on all his classmates than another less popular student. Students may also have different levels of influence on different subgroups, for example, in our empirical results, boys influence boys more and girls influence girls more. This allows us to devise average-maximizing class assignments.
Appendix E Genetic algorithm used to optimize class assignment policy
@@
Algorithm 1 check function
@@
Algorithm 2 swap function
Input: $C$, Set $C$ consists of all students the principal needs to divide into two classrooms
$N_{b}\leftarrow\sum_{s\in C}\mathbbm{1}_{\{s\text{ is boy}\}}$
$N_{g}\leftarrow\sum_{s\in C}\mathbbm{1}_{\{s\text{ is girl}\}}$
$N\leftarrow N_{b}+N_{g}$ Alternatively, $N=\lvert C\rvert$
$j=\operatorname*{arg\,min}_{j}\{N_{j}\}_{j\in\{b,g\}}$
repeat
      $C_{1}\leftarrow C[v]$
      $C_{1}\leftarrow C\backslash C_{1}$ This initialization ensures that $-1\leq\lvert C_{1}\rvert-\lvert C_{2}\rvert\leq 1$
until check$(C_{1},C_{2},N_{b},N_{g},j)=1$ See Algorithm 1 in Appendix E;
L = 150
M = 100
$PH={\rm matrix}(0,L,N)$ Record all the chosen class assignment policies
$FSH={\rm rep}(0,L)$ Record all fitness scores of the chosen policies
for $l$ in $1:L$ do
$b\leftarrow B\sim{\rm Bernoulli}(0.05)$
repeat
       if b = 1 then
            $CP\leftarrow(C_{1}[t_{1}],C_{2}[t_{2}])$ $C_{1}^{\prime},C_{2}^{\prime}\leftarrow{\rm swap}(C_{1},C_{2},CP)$ See Algorithm 2 in Appendix E
until check$(C_{1}^{\prime},C_{2}^{\prime},N_{b},N_{g},j)=1$;
else
      $FS={\rm rep}(0,M)$
      $CS\leftarrow(C_{1}[v_{1}],C_{2}[v_{2}])$ $CS$ denotes M candidate swaps
for $m\in 1:M$ do
       $(C_{1}^{\prime},C_{2}^{\prime})={\rm swap}(C_{1},C_{2},CS_{m})$ $CS_{m}$ is the $m$th candidate swap
      if check$(C_{1}^{\prime},C_{2}^{\prime},N_{b},N_{g},j)=1$ then
            $FS[m]\leftarrow{\rm fit}(C_{1}^{\prime},C_{2}^{\prime})$ ‘fit’ is either (3.5) or (3.6)
      else
            $FS[m]\leftarrow-\infty$
$bp\leftarrow\operatorname*{arg\,max}_{m}FS[m]$
if $FS[bp]>FSH[l-1]$ then
      $CP\leftarrow CS_{bp}$
else
      $CP\leftarrow({\rm NULL},{\rm NULL})$
$(C_{1},C_{2})\leftarrow{\rm swap}(C_{1},C_{2},CP)$
$PH[l,]=(C_{1},C_{2})$
$FSH[l]={\rm fit}(C_{1},C_{2})$
$bp\leftarrow\operatorname*{arg\,max}_{l}FSH[l]$
$C_{1},C_{2}\leftarrow PH[bp,]$
Algorithm 3 GA used to optimize class assignment policy initialization
Appendix F Friendship formation pattern
We show that the friendship formation patterns that we describe in our main text are the same throughout different classroom networks with Figure 5. First, the gender homophily effect is the dominant predictor for friendship formation. Boys almost exclusively make friends with boys and girls make friends with girls. Second, the centrality pattern is clear in the heat maps. Some individual students are extremely influential to other students. Third, students have varying degrees of influence on other students in the same classroom. Even within the same gender, one student is considered by other students as best friends with different probabilities.
Figure 12: Heat maps for two classrooms to demonstrate that PeerNN prediction aligns with known property of social network: (1) gender homophily (2) presence of central nodes (3) heterogeneous popularity across different students
@@ Appendix F Friendship formation pattern
Appendix G Baltagi IV, BVK IV, and full heterogeneous analysis results
See Table 12 and Table 13.
Baltagi IV BVK IV
Peer’s Rank 0.975∗∗∗ 1.075∗∗∗
(0.268) (0.270)
Own Rank 0.997∗∗∗ 0.995∗∗∗
(0.038) (0.038)
Age -0.012∗∗∗ -0.012∗∗∗
(0.001) (0.001)
Sex -0.011 -0.011
(0.017) (0.017)
Father’s education 0.020∗∗∗ 0.020∗∗∗
(0.006) (0.006)
Mother’s education 0.005 0.005
(0.006) (0.006)
Ethnic nationality 0.012 0.012
(0.040) (0.040)
Number of observations 5,860 5,860
Adjusted $R^{2}$ 0.380 0.379
Results from first stage
Peer’s Rank 0.905∗∗∗ 0.905∗∗∗
(0.015) (0.015)
Adjusted $R^{2}$ 0.755 0.755
School FE YES YES
Class RE YES YES
Standard errors in parentheses
$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Table 12: The Peer Effects of Sixth Grade Rank on Cognitive Test Score
@@ Appendix H Class assignment results
We show results for multiple schools, see Figures 13 to 18. From all figures, GA always severely penalizes a few students in pursuit of the highest average peer effect. In contrast, AFGA generates much fairer class assignment policies. In some cases, readers might be able to visually discern that the average peer effect of AFGA is higher than raw assignment, as AFGA tends to have lighter color plots.
@@
Gender Father’s education Types of Hukou
Male Female Low High Rural Others
Peer’s Rank 1.000 0.946 1.255∗∗∗ 0.565 1.215∗∗ 1.143
(0.581) (0.592) (0.582) (0.687) (0.619) (0.650)
Own Rank 0.916∗∗∗ 1.117∗∗∗ 0.990∗∗∗ 1.037∗∗∗ 0.984∗∗∗ 1.008∗∗∗
(0.054) (0.053) (0.053) (0.056) (0.055) (0.053)
Age -0.013∗∗∗ -0.011∗∗∗ -0.012∗∗∗ -0.011∗∗∗ -0.012∗∗∗ -0.012∗∗∗
(0.002) (0.002) (0.002) (0.002) (0.002) (0.002)
Sex 0.006 -0.032 0.002 -0.030
(0.025) (0.024) (0.026) (0.023)
Father’s education 0.016 0.024∗∗∗ 0.011 0.019∗∗
(0.009) (0.008) (0.010) (0.008)
Mother’s education 0.007 0.003 0.008 0.008 -0.003 0.006
(0.009) (0.008) (0.011) (0.007) (0.011) (0.008)
Ethnic nationality 0.121∗∗ -0.062 0.033 -0.002 0.100 -0.026
(0.062) (0.051) (0.060) (0.052) (0.067) (0.049)
Number of observations 3,048 2,812 3,147 2,713 2,845 3,015
- Log Likelihood 3,244.771 2,611.928 3,283.262 2,561.519 2,953.243 2,924.823
School FE YES YES YES YES YES YES
Class RE YES YES YES YES YES YES
Standard errors in parentheses
$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Table 13: Heterogeneous Peer Effects of Sixth Grade Rank on Cognitive Test Score
Appendix I Balance tests for class random assignment
To assess whether class assignment is based on students’ academic outcome, we implement a balance check regression examining the relationship between students’ own characteristics and their classmates’ 6th-grade class quantile (leave-out mean). Specifically, we regress the class quantile (excluding the individual student) on the student’s own characteristics to test whether these characteristics predict classmates’ academic ranks. The results, shown in Table 14, indicate that after controlling for school fixed effects, all coefficients are statistically insignificant except for Hukou type and age. The coefficient on Hukou type is marginally significant at the 5% level. This is reasonably expected, as students within a school are typically drawn from the same geographic area and therefore tend to share the same Hukou type (rural or urban). As a result, a student’s own Hukou status may weakly predict that of their classmates, introducing a slight correlation with the classmates’ class quantile. The coefficient on age is significant at the 1% level. However, it is unlikely that schools assign students to classes based on age, particularly since grade cohorts are generally age-homogeneous. Even though age is measured in months, students within the same grade and school are expected to have similar ages. This is supported by Figure 19, which displays the distribution of within-school age variance. The red dashed line marks the full-sample variance of age, and the figure shows that most schools have low within-school variance, suggesting little variation in student age within schools. Therefore, these results support the assumption that class assignment within schools is random.
Peer’s Rank
Age -0.0005∗∗∗
Peer’s age -0.022∗∗∗
Sex -0.001
Hukou -0.002
Father’s education 0.0001
Mother’s education 0.0001
Ethnic nationality 0.0003
Number of observations 5,860
Adjusted $R^{2}$ 0.192
School FE YES
Class RE NO
Standard errors in parentheses
$p<0.10$, ∗∗ $p<0.05$, ∗∗∗ $p<0.01$
Table 14: Balance test regression based on class rank
Figure 19: Density of Within-school age distribution
Figure 20: Distribution of P-values from Pearson $\chi^{2}$ Test